# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 10 Densification — Phase 3
By Lemma 5.5, we have $|v_{1}(i)|\leq(1+2p+2/n)/\sqrt{n}$ for every $i\in[n]$, thus
$1$ $=\sum_{i=1}^{n}v_{1}(i)^{2}\leq|U|\frac{(1+2p+2/n)^{2}}{n}+(n-|U|)\frac{(1-8p)^{2}}{n}$
$\leq\frac{1}{n}\left(|U|(1+8p)+(n-|U|)(1-8p)\right)=(1-8p)+\frac{16p|U|}{n}.$
From this, we get $|U|\geq n/2$. ∎
We now evaluate the terms of $\mathds{1}_{U}^{T}D\mathds{1}_{U}$, starting with the main term $\mathds{1}_{U}^{T}B^{\circ 3}\mathds{1}_{U}$.
@@
$\mathds{1}^{T}_{U}B^{\circ 3}\mathds{1}_{U}\leq 10^{6}|U|^{2}p^{3}-e(\overline{G}[U])/4$.
@@
We have
If $i,j\in I$, then
$1-\lambda_{1}v_{1}(i)v_{1}(j)\leq 1-(1-p)(n-1)\cdot\left(\frac{1-8p}{\sqrt{n}}\right)^{2}\leq 100p,$
and thus $\lambda_{1}v_{1}(i)v_{1}(j)>\frac{1}{2}$. Therefore,
$\mathds{1}^{T}_{U}B^{\circ 3}\mathds{1}_{U}=\sum_{i,j\in I,i\sim j}(1-\lambda_{1}v_{1}(i)v_{1}(j))^{3}-\sum_{i,j\in U,i\not\sim j}(\lambda_{1}v_{1}(i)v_{1}(j))^{3}\leq 100^{3}|U|^{2}p^{3}-2e(\overline{G}[U])/8.$
Claim 10.4.
We have
$\mathds{1}^{T}_{U}(3|\lambda_{n}|B\circ B\circ I+3|\lambda_{n}|B\circ I+|\lambda_{n}|^{3}I)\mathds{1}_{U}\leq O(n^{1+3s}).$
Proof.
We have $B_{i,i}=-\lambda_{1}v_{1}(i)^{2}$ for every $i\in[n]$. By Lemma 5.5, we have $|B_{i,i}|\leq 2$. Then the claim follows immediately as the expression above evaluates to
$\sum_{i\in U}(3|\lambda_{n}|B_{i,i}^{2}+3|\lambda_{n}|^{2}B_{i,i}+|\lambda_{n}|^{3})\leq O(n|\lambda_{n}|^{3}).$
In conclusion, we proved that
$0\leq\mathds{1}^{T}_{U}D\mathds{1}_{U}\leq 10^{6}|U|^{2}p^{3}-e(\overline{G}[U])/4+O(n^{1+3s}).$
Hence, we must have $e(\overline{G}[U])\leq 10^{7}|U|^{2}p^{3}/2+O(n^{1+3s})$, which shows that the edge density of $G[U]$ is at least $1-10^{7}p^{3}-O(n^{3s-1})$. ∎
@@
Let $0<s<1/3$, and let $n$ be sufficiently large with respect to $s$. Let $G$ be an $n$-vertex graph with edge density at least $1-10^{-5}$. If $|\lambda_{n}|<n^{s}$, then $G$ contains an induced subgraph of size $n^{1-o(1)}$ with edge density at least $1-n^{3s-1+o(1)}$.
Proof.
Let $s_{0}$ be such that $s<s_{0}<1/3$, and let 1p1-p be the edge density of $G$, $p<10^{-5}$. Let $G_{0}=G$, and define the sequence of induced subgraphs $G_{0}\supset G_{1}\supset$... with increasing density as follows. If $G_{i}$ is already defined with $n_{i}$ vertices and edge density $1-p_{i}$, then stop if either $p_{i}<n_{i}^{3s_{0}-1}$, or $n_{i}<n^{s/s_{0}}$. Otherwise, as $n_{i}>n^{s/s_{0}}$ and the smallest eigenvalue of $G_{i}$ is at least $\lambda_{n}>-n^{s}>-n_{i}^{s_{0}}$, we can apply Lemma 10.1 to find an induced subgraph $G_{i+1}$ of $G_{i}$ on at least $n_{i}/2$ vertices of edge density at least $1-10^{7}p_{i}^{3}-O(n_{i}^{3s_{0}-1})\geq 1-\max\{10^{8}p_{i}^{3},O(n_{i}^{3s_{0}-1})\}$.
Let $L$ be the last index $i$ for which $G_{i}$ is defined. Then $n_{L}\geq n2^{-L}$. On the other hand $p_{L-1}>n_{L-1}^{3s_{0}-1}$, and so for every $i\leq L-2$, we have $p_{i+1}\leq 10^{8}p_{i}^{3}$. This show that
$p_{L-1}<(10^{8})^{3^{L-2}+3^{L-3}+\dots+1}p^{3^{L-1}}<(10^{4}p)^{3^{L-1}}<10^{-3^{L-1}}.$
Hence, we must have stopped because $p_{L}<n_{i}^{3s_{0}-1}$, which happens for some $L=O(\log\log n)$. But then $n_{L}\geq n4^{-O(\log\log n)}=n^{1-o(1)}$. We thus get that $G_{L}$ is a subgraph of $G$ of size $n^{1-o(1)}$ with edge density at least $1-n^{3s_{0}-1+o(1)}$. As we may choose $s_{0}$ arbitrarily close to $s$, this finishes the proof. ∎
Now we establish the MaxCut version of Lemma 10.1. While the core idea remains the same, the proof becomes more involved. Therefore, we give a detailed proof.
Lemma 10.6.
Let $0<s<1/4$ and $0<\alpha<\min\{\frac{1}{12}-\frac{s}{6},\frac{1}{6}-\frac{2s}{3}\}$, and let $n$ be sufficiently large with respect to $s,\alpha$. Let $G$ be an $n$-vertex graph with edge density 1p1-p, where $n^{-\alpha}\leq p<10^{-5}$. Assume that $\operatorname{surp}^{*}(G)<n^{1+s}$. Then $G$ contains an $n/4$-vertex induced subgraph of edge density at least $1-10^{8}p^{3}$.
Proof.
Let $A$ be the adjacency matrix of $G$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and corresponding orthonormal basis of eigenvectors $v_{1},\dots,v_{n}$. Define $B=A-\lambda_{1}v_{1}v_{1}^{T}$ and $E=\sum_{\lambda_{i}<0}|\lambda_{i}|v_{i}v_{i}^{T}$. Then, the matrices $E$ and $B+E=\sum_{\lambda_{i}>0,i\neq 1}\lambda_{i}v_{i}v_{i}^{T}$ are positive semidefinite.
The key idea of the proof is to consider the following triple Hadamard product:
As $B+E$ is positive semidefinite, so is $D$ by the Schur product theorem. Then, using the matrix $E$ and the principal eigenvector $v_{1}$, we identify a set of well-behaved vertices $U$, and evaluate the product
$0\leq\mathds{1}_{U}^{T}D\mathds{1}_{U}=\mathds{1}_{U}^{T}B^{\circ 3}\mathds{1}_{U}+\mathds{1}_{U}^{T}\big{(}3B\circ B\circ E+3B\circ E\circ E+E^{\circ 3}\big{)}\mathds{1}_{U}.$
By carefully analyzing the terms of this product, we will conclude that the graph $\overline{G}[U]$ must be much sparser than $G$.
We now give the details. First, observe that
$\operatorname{tr}(E)=\sum_{\lambda_{i}<0}|\lambda_{i}|\leq\operatorname{surp}^{*}(G)\leq n^{1+s}$
by (i) in Lemma 5.3, and
$\|E\|_{F}^{2}=\sum_{\lambda_{i}<0}\lambda_{i}^{2}\leq O(\sqrt{\Delta+1})\operatorname{surp}^{*}(G)\leq O(\sqrt{\Delta+1})n^{1+s}\leq O(n^{3/2+s})$
by (ii) in Lemma 5.3. Let $U$ be the set of vertices $i\in[n]$ that satisfy $E_{i,i}\leq 4n^{s}$ and $v_{1}(i)\geq(1-8p)/\sqrt{n}$.
Claim 10.7.
$|U|\geq n/4$.
Proof.
Let $U_{0}$ be the set of vertices $i\in[n]$ such that $v_{1}(i)\geq(1-8p)/\sqrt{n}.$ Then using Lemma 5.5,
$1$ $=\sum_{i=1}^{n}v_{1}(i)^{2}\leq|U_{0}|\frac{(1+2p+2/n)^{2}}{n}+(n-|U_{0}|)\frac{(1-8p)^{2}}{n}$
$\leq\frac{1}{n}\left(|U_{0}|(1+8p)+(n-|U_{0}|)(1-8p)\right)=(1-8p)+\frac{16p|U_{0}|}{n}.$
From this, we get $|U_{0}|\geq n/2$. Now $U$ is those set of vertices $i\in U_{0}$ that satisfy $E_{i,i}\leq 4n^{s}$. As $\operatorname{tr}(E_{i,i})\leq n^{1+s}$, the number of vertices such that $E_{i,i}>4n^{s}$ is at most $n/4$, giving the desired bound $|U|\geq n/4$. ∎
[…]
$\mathds{1}^{T}_{U}B^{\circ 3}\mathds{1}_{U}\leq 10^{6}|U|^{2}p^{3}-e(\overline{G}[U])/4.$
So we consider the rest of the terms.
Claim 10.8.
We have
$\mathds{1}^{T}_{U}(3B\circ B\circ E+3B\circ E\circ E+E^{\circ 3})\mathds{1}_{U}\leq O(n^{7/4+s/2}+n^{3/2+2s}).$
Proof.
We bound each summand of the error term independently. Firstly, note that every entry of $B$ is between $1$ and 2-2, as $0\leq\lambda_{1}v_{1}(i)v_{2}(j)\leq n\left(\frac{1+2p+2/n}{\sqrt{n}}\right)^{2}\leq 2$. Therefore, we have
$\mathds{1}^{T}_{U}(3B\circ B\circ E)\mathds{1}_{U}\leq 12\sum_{i,j\in U}|E_{i,j}|\leq 12|U|\sqrt{\sum_{i,j\in U}E_{i,j}^{2}}\leq 12n\|E\|_{F}\leq O(n^{7/4+s/2}).$
Here, the second inequality holds by the inequality between the arithmetic and quadratic mean.
To bound the second summand, we again use that entries of $B$ are bounded by $2$ in absolute value, and so
$\mathds{1}^{T}_{U}(3B\circ E\circ E)\mathds{1}_{U}\leq 6\sum_{i,j\in U}E_{i,j}^{2}\leq 6\|E\|_{F}^{2}\leq O(n^{3/2+s}).$
Finally, in bounding the last term we use that $E_{i,i}\leq n^{s}$ for all $i\in U$. In particular, since $E$ is a positive definite matrix, this implies that $|E_{i,j}|\leq n^{s}$ for all $i,j\in U$. So,
$\mathds{1}^{T}_{U}E^{\circ 3}\mathds{1}_{U}\leq\sum_{i,j\in U}|E_{i,j}|^{3}\leq\max_{i,j\in U}|E_{i,j}|\cdot\sum_{i,j\in U}|E_{i,j}|^{2}\leq n^{s}\cdot\|E\|_{F}^{2}\leq O(n^{3/2+2s}).$
The conclusion now follows by summing up the bounds obtained on each of the error terms above. ∎
In conclusion, we proved that
$0\leq\mathds{1}^{T}_{U}D\mathds{1}_{U}\leq 10^{6}|U|^{2}p^{3}-e(\overline{G}[U])/4+O(n^{7/4+s/2}+n^{3/2+2s})\leq 10^{7}|U|^{2}p^{3}-e(\overline{G}[U])/4,$
where in the last inequality we used our bounds on $p$ and $s$, and that $n$ is sufficiently large. Hence, we must have $e(\overline{G}[U])\leq 10^{8}|U|^{2}p^{3}/2$, which shows that the edge density of $G[U]$ is at least $(1-10^{8}p^{3})$. ∎
Lemma 10.9.
Let $0<s<1/4$ and $0<\alpha<\min\{\frac{1}{12}-\frac{s}{6},\frac{1}{6}-\frac{2s}{3}\}$, and let $n$ be sufficiently large with respect to $s,\alpha$. Let $G$ be an $n$-vertex graph with edge density at least $1-10^{-5}$. If $\operatorname{surp}^{*}(G)<n^{1+s}$, then $G$ contains an induced subgraph of size $n^{1-o(1)}$ with edge density at least $1-n^{-\alpha}$.
Proof.
The proof of this is almost identical to the proof of Lemma 10.5, but instead on using Lemma 10.1, we use Lemma 10.6. Then, we just need to adjust the parameters $s$ and $\alpha$ slightly. We skip further details. ∎
11 Densification — Phase 4
In this section, we show that graphs with large smallest eigenvalue (or small surplus) of edge density $1-n^{-\varepsilon}$ contain subgraphs of edge density $1-n^{-1+\varepsilon}$ on almost the same number of vertices.
@@ 11 Densification — Phase 4
In order to prove this, it is more convenient to work with the complement $\overline{G}$ of $G$. Let $\lambda_{1}\geq\dots\geq\lambda_{n}$ be the eigenvalues of $G$, and let $\mu_{1}\geq\dots\geq\mu_{n}$ be the eigenvalues of $\overline{G}$. Unfortunately, as $G$ is not necessarily regular, there is no simple formula to express $\mu_{i}$ in terms of $\lambda_{1},\dots,\lambda_{n}$. However, we can use Weyl’s inequality to establish the following interlacing property.
@@
where $\lambda_{1}(X)\geq\dots\geq\lambda_{n}(X)$ denote the eigenvalues of a matrix $X$. Let $A$ be the adjacency matrix of $G$, and let $B$ be the adjacency matrix of $\overline{G}$. Then B=JIAB=J-I-A. Let X=AX=-A and Y=JIY=J-I, then $\lambda_{i}(X)=-\lambda_{n+1-i}$, $\lambda_{1}(Y)=n-1$, $\lambda_{i}(X)=-1$ for $i=2,\dots,n$, and $\lambda_{i}(X+Y)=\mu_{i}$. Hence, applying the above inequality with $j=2$, we get
$\mu_{i+1}\leq-\lambda_{n+1-i}-1.$
The next lemma provides a bound on the surplus of very dense graphs. The result and its proof are similar to the proof of Lemma 5.9 of Räty, Sudakov and Tomon [64] for the complementary quantity called as positive discrepancy. Say that a graph $G$ is $C$-balanced if $\Delta(G)\leq Cd(G)$.
Lemma 11.2.
Let $G$ be an $n$-vertex graph of density (1p)(1-p) such that the complement of $G$ is $C$-balanced, and $p<0.001C^{-2}$. Then
$\operatorname{surp}^{*}(G)\geq\Omega\left(\min\left\{\frac{n}{C^{3}p},C^{-1}p^{1/2}n^{3/2}\right\}\right).$
Proof.
Let $A$ be the adjacency matrix of $G$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$, and let $B$ be the adjacency matrix of $\overline{G}$ with eigenvalues $\mu_{1}\geq\dots\geq\mu_{n}$.
Let $\bar{\Delta}$ be the maximum degree of $\overline{G}$, so $\mu_{1}\leq\bar{\Delta}\leq Cpn$. We may assume that $p>0$ and thus $\bar{\Delta}\geq 1$, otherwise the statement is trivial. For $k=1,2,3$, set
@@
Lemma 11.1 applied with $i\geq 2$ shows that whenever $\mu_{i}\geq 0$ we also have $\lambda_{n+1-i}\leq-\mu_{i}-1<0$. Combined with Lemma 5.3, this shows that
$\operatorname{surp}^{*}(G)$ $\geq\sum_{\lambda_{i}<0}|\lambda_{i}|\geq\sum_{i\neq 1,\mu_{i}>0}\mu_{i}=P_{1},$
$\operatorname{surp}^{*}(G)$ $\geq\Omega\left(\bar{\Delta}^{-1/2}\sum_{\lambda_{i}<0}|\lambda_{i}|^{2}\right)\geq\Omega\left({\bar{\Delta}^{-1/2}}\sum_{i\neq 1,\mu_{i}>0}\mu_{i}^{2}\right)=\Omega\left(\bar{\Delta}^{-1/2}P_{2}\right),$
$\operatorname{surp}^{*}(G)$ $\geq\Omega\left(\frac{1}{\bar{\Delta}}\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}\right)\geq\Omega\left(\frac{1}{\bar{\Delta}}\sum_{i\neq 1,\mu_{i}>0}\mu_{i}^{3}\right)=\Omega\left(\frac{1}{\bar{\Delta}}P_{3}\right).$

### v2

@@ 9.1 Finding balanced subgraphs
Let $U\subseteq V(G)$ be any subset of size at least (11/4C)n(1-1/4C)n. Since $G$ is $C$-balanced, the number of edges with an endpoint outside $U$ is at most $\frac{n}{4C}\cdot\Delta(G)\leq\frac{n}{4C}\cdot Cp(n-1)=n(n-1)p/4$. So $e(G[U])\geq n(n-1)p/4$ and the density of $G[U]$ is at least $\frac{n(n-1)p/4}{|U|(|U|-1)/2}\geq p/2$. ∎
9.2 The smallest eigenvalue and surplus of dense balanced graphs
In this section, we prove that balanced graphs with a sufficiently high density 1p1-p must satisfy $|\lambda_{n}|=\Omega((pn)^{1/3})$ and $\operatorname{surp}^{*}(G)=\Omega((pn)^{1/4})$. Both bounds are effective as long as $p$ is not too small. Then, in the next subsection, we show complementary bounds for small $p$.
First, we consider the smallest eigenvalue.
Lemma 9.4.
Let $G$ be an $n$-vertex graph with edge density 1p1-p, whose complement $\overline{G}$ is $C$-balanced. If $C^{2}p\leq 1/100$, then
$|\lambda_{n}|=\Omega\big((pn)^{1/3}\big).$
Proof.
We may assume $p\geq 1/n$ as otherwise $|\lambda_{n}|\geq 1\geq(pn)^{1/3}$. Let $A$ be the adjacency matrix of $G$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and corresponding orthonormal basis of eigenvectors $v_{1},\dots,v_{n}$. Set $B=A-\lambda_{1}v_{1}v_{1}^{T}$. The key idea of the proof is to consider the following triple Hadamard product:
$D=\big(B+|\lambda_{n}|I\big)^{\circ 3}=B^{\circ 3}+3|\lambda_{n}|B\circ B\circ I+3|\lambda_{n}|^{2}B\circ I+|\lambda_{n}|^{3}I.$
As $B+|\lambda_{n}|I=|\lambda_{n}|v_{1}v_{1}^{T}+\sum_{i=2}^{n}(\lambda_{i}+|\lambda_{n}|)v_{i}v_{i}^{T}$, we have that $B+|\lambda_{n}|I$ is positive semidefinite. Therefore, $D$ is also positive semidefinite by the Schur product theorem (Theorem˜4.1).
Write $\bar{\Delta}$ for the maximum degree $\overline{G}$; then $\bar{\Delta}\leq Cpn$. By Lemma˜5.5, for every vertex $i\in V(G)$,
$0\leq\frac{1-3Cp}{\sqrt{n}}\leq\frac{1-3\bar{\Delta}/n}{\sqrt{n}}\leq v_{1}(i)\leq\frac{1+2p+2/n}{\sqrt{n}}\leq\frac{1+4p}{\sqrt{n}}$ (8)
We now evaluate the terms of $\mathds{1}^{T}D\mathds{1}$, where $\mathds{1}$ is the all-ones vector.
For the main term $\mathds{1}^{T}B^{\circ 3}\mathds{1}$, note that
Using that $\lambda_{1}\geq d(G)=(1-p)(n-1),C\geq 1,p\geq 1/n$ and $Cp\leq C^{2}p\leq 1/100$, we have for every $i,j\in V(G)$ that
$1-\lambda_{1}v_{1}(i)v_{1}(j)\leq 1-(1-p)(n-1)\cdot\frac{(1-3Cp)^{2}}{n}\leq p+6Cp+\frac{1}{n}\leq 8Cp\leq 1/2.$
This further shows $\lambda_{1}v_{1}(i)v_{1}(j)\geq\frac{1}{2}$. Hence, using that $C^{3}p^{2}\leq(C^{2}p)^{2}\leq 10^{-4}$, we get
$\mathds{1}^{T}B^{\circ 3}\mathds{1}$ $=\sum_{i\sim j}(1-\lambda_{1}v_{1}(i)v_{1}(j))^{3}-\sum_{i\not\sim j}(\lambda_{1}v_{1}(i)v_{1}(j))^{3}$
[…]
For the other terms, we observe that $B_{i,i}=-\lambda_{1}v_{1}(i)^{2}\in[-2,0]$ using (8). Hence,
$\mathds{1}^{T}\big(3|\lambda_{n}|B\circ B\circ I+3|\lambda_{n}|^{2}B\circ I+|\lambda_{n}|^{3}I\big)\mathds{1}=O\big(n|\lambda_{n}|^{3}\big).$
In conclusion, we showed that
$0\leq\mathds{1}^{T}D\mathds{1}\leq-\frac{pn^{2}}{16}+O\big(n|\lambda_{n}|^{3}\big).$
This gives $|\lambda_{n}|=\Omega\big((pn)^{1/3}\big)$, as desired. ∎
Now we consider the surplus.
Lemma 9.5.
Let $G$ be an $n$-vertex graph with edge density 1p1-p, whose complement $\overline{G}$ is $C$-balanced. If $C^{2}p\leq 1/100$, then
$\operatorname{surp}^{*}(G)=\Omega\big(C^{-3/4}n(pn)^{1/4}\big).$
Proof.
Set $S:=\operatorname{surp}^{*}(G)$ and assume for contradiction that $S=o\big(C^{-3/4}n(pn)^{1/4}\big)$. Since $p\leq C^{2}p\leq 1/100$, there are $\Omega(n)$ non-isolated vertices in $G$. Then, a classical result of Erdős, Gyárfás, and Kohayakawa [41] asserts $\operatorname{surp}(G)=\Omega(n)$, so $\operatorname{surp}^{*}(G)\geq\operatorname{surp}(G)=\Omega(n)$. Hence, $S=o\big(C^{-3/4}n(pn)^{1/4}\big)$ implies $pn=\omega(1)$ and $C=o\big((pn)^{1/3})\big)$.
Let $A$ be the adjacency matrix of $G$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and corresponding orthonormal basis of eigenvectors $v_{1},\dots,v_{n}$. Define $B=A-\lambda_{1}v_{1}v_{1}^{T}$ and $E=\sum_{\lambda_{i}<0}|\lambda_{i}|v_{i}v_{i}^{T}$. Then, the matrices $E$ and $B+E=\sum_{\lambda_{i}>0,i\neq 1}\lambda_{i}v_{i}v_{i}^{T}$ are positive semidefinite. Consider
As $B+E$ is positive semidefinite, so is $D$ by the Schur product theorem (Theorem˜4.1). Next, we identify a set of well-behaved vertices $U$, and carefully evaluate the product
$0\leq\mathds{1}_{U}^{T}\,D\,\mathds{1}_{U}=\mathds{1}_{U}^{T}\,B^{\circ 3}\,\mathds{1}_{U}+3\cdot\mathds{1}_{U}^{T}\big(B\circ B\circ E\big)\mathds{1}_{U}+3\cdot\mathds{1}_{U}^{T}\big(B\circ E\circ E\big)\mathds{1}_{U}+\mathds{1}_{U}^{T}\,E^{\circ 3}\,\mathds{1}_{U}.$ (9)
Let $\bar{d},\bar{\Delta}$ be the average degree and the maximum degree of $\overline{G}$, respectively, so we have $\bar{\Delta}\leq C\bar{d}$. Let $U$ be the set of vertices $i\in[n]$ such that $E_{i,i}\leq\frac{4CS}{n}$. Recall from Lemma˜5.3 (i) that
$\operatorname{tr}(E)=\sum_{\lambda_{i}<0}|\lambda_{i}|\leq\operatorname{surp}^{*}(G)=S.$
So at most $n/4C$ vertices $i\in[n]$ have $E_{i,i}>\frac{4CS}{n}$. This means $|U|\geq(1-\frac{1}{4C})n$. By ˜9.3, the density of $\overline{G}[U]$ is at least $p/2$. Moreover, Lemma˜5.3 (ii) and our assumption that $S=o\big(n\bar{d}^{\,1/4}C^{-3/4}\big)$ imply
$\|E\|_{F}^{2}=\sum_{\lambda_{i}<0}\lambda_{i}^{2}=O\Big({\bar{\Delta}}^{1/2}\operatorname{surp}^{*}(G)\Big)=O\big(C^{1/2}\bar{d}^{\,1/2}S\big)=o\big(n\bar{d}^{\,3/4}C^{-1/4}\big).$ (10)
[…]
$\frac{1-3Cp}{\sqrt{n}}\leq v_{1}(i)\leq\frac{1+4p}{\sqrt{n}}\;\;\text{ and }\;\;1-\lambda_{1}v_{1}(i)v_{1}(j)\leq 8Cp.$
Using that $p\leq C^{2}p\leq 1/100$, for every $ij\in E(G)$, we have
$1\leq-9p\leq 1-(1+4p)^{2}\leq B_{i,j}=1-\lambda_{1}v_{1}(i)v_{1}(j)\leq 8Cp\leq 1,$
while for any $ij\not\in E(G)$, using that $p\leq Cp\leq C^{2}p\leq 1/100$, we have
$2\leq-(1+4p)^{2}\leq B_{i,j}=-\lambda_{1}v_{1}(i)v_{1}(j)\leq 8Cp-1\leq-\frac{1}{2}.$
Since $C^{3}p^{2}\leq(C^{2}p)^{2}\leq 10^{-4}$,
$\mathds{1}^{T}_{U}B^{\circ 3}\mathds{1}_{U}$ $=\sum_{i,j\in U,i\sim j}(1-\lambda_{1}v_{1}(i)v_{1}(j))^{3}-\sum_{i,j\in U,i\not\sim j}(\lambda_{1}v_{1}(i)v_{1}(j))^{3}$
[…]
In addition, $|U|\geq(1-\frac{1}{4C})n=\Omega(n)$, so $\mathds{1}^{T}_{U}B^{\circ 3}\mathds{1}_{U}=-\Omega(pn^{2})=-\Omega(n\bar{d})$.
We complete the proof by the following claim bounding the rest of the terms of (9).
@@ 9.2 The smallest eigenvalue and surplus of dense balanced graphs
  • (1)
    $\mathds{1}_{U}^{T}\big(B\circ B\circ E\big)\mathds{1}_{U}=o(n\bar{d}),$
  • (2)
    $\mathds{1}_{U}^{T}\big(B\circ E\circ E\big)\mathds{1}_{U}=o(n\bar{d}),$
  • (3)
    $\mathds{1}_{U}^{T}\,E^{\circ 3}\,\mathds{1}_{U}=o(n\bar{d}).$
@@
  • (1)
    Write $X$ for the set of pairs $(i,j)\in U^{2}$ where $ij\in E(G)$ and $Y=(U\times U)\setminus X$. As discussed above, $|B_{i,j}|=O(Cp)$ for all $(i,j)\in X$. So, by the Cauchy–Schwartz inequality,
    $\sum_{(i,j)\in X}B_{i,j}^{2}E_{i,j}=O(C^{2}p^{2})\sum_{(i,j)\in X}|E_{i,j}|=O(C^{2}p^{2})\cdot n\Big(\sum_{(i,j)\in X}|E_{i,j}|^{2}\Big)^{1/2}=O\big(C^{2}p^{2}n\|E\|_{F}\big).$
    Using that $C^{2}p^{2}\leq(C^{2}p)^{3/2}p^{1/2}\leq p^{1/2}$, we see that the above is at most $O\big(p^{1/2}n\|E\|_{F}\big)$. Moreover, $|Y|\leq n^{2}p+n\leq 2n^{2}p$ and $|B_{i,j}|\leq 2$ for all $(i,j)\in Y$. Again, by the Cauchy–Schwarz inequality,
    $\sum_{(i,j)\in Y}B_{i,j}^{2}E_{i,j}\leq 4\sum_{(i,j)\in Y}|E_{i,j}|\leq 4\cdot\Big(|Y|\sum_{(i,j)\in Y}|E_{i,j}|^{2}\Big)^{1/2}=O\big(p^{1/2}n\|E\|_{F}\big).$
    Altogether, we have
    $\mathds{1}_{U}^{T}\big(B\circ B\circ E\big)\mathds{1}_{U}=\sum_{(i,j)\in X}B_{i,j}^{2}E_{i,j}+\sum_{(i,j)\in Y}B_{i,j}^{2}E_{i,j}=O\big(p^{1/2}n\|E\|_{F}\big)=O\big((n\bar{d})^{1/2}\|E\|_{F}\big).$
    By (10), $\|E\|_{F}^{2}=o(n\bar{d}^{\,3/4}C^{-1/4})=o(n\bar{d}^{\,3/4})$, so
    $\mathds{1}_{U}^{T}\big(B\circ B\circ E\big)\mathds{1}_{U}=o\big((n\bar{d})^{1/2}\cdot n^{1/2}\bar{d}^{\,3/8})\big)=o(n\bar{d}).$
  • (2)
    As $|B_{i,j}|\leq 2$ holds for all $i,j\in U$ and $\|E\|_{F}^{2}=o\big(n\bar{d}^{\,3/4}\big)$, we have
    $\mathds{1}_{U}^{T}\big(B\circ E\circ E\big)\mathds{1}_{U}=\sum_{i,j\in U}B_{i,j}E_{i,j}^{2}\leq 2\|E\|_{F}^{2}=o\big(n\bar{d}^{\,3/4}\big).$
  • (3)
    Recall that $E_{i,i}\leq 4CS/n=o\big((C\bar{d})^{1/4}\big)$ for all $i\in U$. As $E$ is positive semidefinite, this implies that $|E_{i,j}|\leq 4CS/n=o\big((C\bar{d})^{1/4}\big)$ for all $i,j\in U$. This, together with (10), completes the proof as
    $\mathds{1}_{U}^{T}\,E^{\circ 3}\,\mathds{1}_{U}=\sum_{i,j\in U}E_{i,j}^{3}=\max_{i,j\in U}|E_{i,j}|\cdot\sum_{i,j\in U}E_{i,j}^{2}=o\big((C\bar{d})^{1/4}\big)\cdot\|E\|_{F}^{2}=o(n\bar{d}).\hfill\qed$
With these estimates, we can rewrite (9) as $0\leq\mathds{1}^{T}_{U}\,D\,\mathds{1}_{U}\leq-\Omega(n\bar{d})+o(n\bar{d})$; a contradiction. ∎
9.3 Further bounds on balanced graphs
In this subsection, we prove that if $G$ is a balanced graph of sufficiently large density 1p1-p, then $|\lambda_{n}|\geq\frac{\operatorname{surp}^{*}(G)}{n}=\Omega(\min\{p^{-1},(pn)^{1/2}\})$, where the first inequality holds by ˜5.7. Note that this beats the lower bound for $|\lambda_{n}|$ from the previous section when $p\ll n^{-1/4}$, and beats the lower bound for the surplus when $p\ll n^{-1/5}$.
@@ 9.3 Further bounds on balanced graphs
Let $G$ be an $n$-vertex graph with edge density 1p1-p, whose complement $\overline{G}$ is $C$-balanced. If $C^{2}p\leq 1/100$, then
$\operatorname{surp}^{*}(G)=\Omega\Big(C^{-3}\cdot\min\{np^{-1},n(pn)^{1/2}\}\Big).$
In particular,
$|\lambda_{n}|=\Omega\Big(C^{-3}\cdot\min\{p^{-1},(pn)^{1/2}\}\Big).$
@@
To establish these lower bounds, it is more convenient to work with the complement of $G$. Unfortunately, as $G$ is not necessarily regular, there is no simple formula to express the eigenvalues of $\overline{G}$ in terms of those of $G$. However, we can use Weyl’s inequality to establish the following inequality.
@@
where $\lambda_{1}(X)\geq\dots\geq\lambda_{n}(X)$ denote the eigenvalues of a matrix $X$. Let $A$ be the adjacency matrix of $G$, and then A+JI-A+J-I is the adjacency matrix of $\overline{G}$. Set X=AX=-A and Y=JIY=J-I. We have $\lambda_{i}(X)=-\lambda_{n+1-i}$, $\lambda_{1}(Y)=n-1$, $\lambda_{i}(Y)=-1$ for $i=2,\dots,n$, and $\lambda_{i}(X+Y)=\mu_{i}$. Hence, applying the above inequality with $j=2$, we get for $i=1,2,\dots,n-1$,
$\mu_{i+1}=\lambda_{i+1}(X+Y)\leq\lambda_{i}(X)+\lambda_{2}(Y)=-\lambda_{n+1-i}-1.\hfill\qed$
Proof of Lemma˜9.7.
We focus on the lower bounds for $\operatorname{surp}^{*}(G)$ since that of $\lambda_{n}$ follows from ˜5.7, i.e. $|\lambda_{n}|n\geq\operatorname{surp}^{*}(G)$.
Let $A$ be the adjacency matrix of $G$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$, and let $B$ be the adjacency matrix of $\overline{G}$ with eigenvalues $\mu_{1}\geq\dots\geq\mu_{n}$. Write $\bar{\Delta}$ for the maximum degree of $\overline{G}$, so $\mu_{1}\leq\bar{\Delta}\leq Cpn$. We may assume that $p>0$ and thus $\bar{\Delta}\geq 1$, as otherwise the statement is trivial. For $k=1,2,3$, set
@@
By Lemma˜9.8, whenever $\mu_{i+1}\geq 0$, we also have $\lambda_{n+1-i}\leq-\mu_{i+1}-1<0$. Combined with Lemma˜5.3, this shows that
$\operatorname{surp}^{*}(G)$ $\geq\sum_{\lambda_{i}<0}|\lambda_{i}|\geq\sum_{i\neq 1,\mu_{i}>0}\mu_{i}=P_{1},$ (11)
$\operatorname{surp}^{*}(G)$ $=\Omega\bigg(\bar{\Delta}^{-1/2}\sum_{\lambda_{i}<0}|\lambda_{i}|^{2}\bigg)=\Omega\bigg({\bar{\Delta}^{-1/2}}\sum_{i\neq 1,\mu_{i}>0}\mu_{i}^{2}\bigg)=\Omega\Big(\bar{\Delta}^{-1/2}P_{2}\Big),$ (12)
$\operatorname{surp}^{*}(G)$ $=\Omega\bigg(\bar{\Delta}^{-1}\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}\bigg)=\Omega\bigg(\bar{\Delta}^{-1}\sum_{i\neq 1,\mu_{i}>0}\mu_{i}^{3}\bigg)=\Omega\Big(\bar{\Delta}^{-1}P_{3}\Big).$ (13)